Solve by clearing fractions:
7/3-5/2x=8+2x

Answer by stanbon(75887)

7/3-5/2x=8+2x
----------------
Multiply thru by 6 to get:
14 - 15x = 8+2x
----
17x = 6
x = 6/17
===================
Cheers,
Stan H.